# Certain Inequalities for the generalized polar derivative of a polynomial

### v1

@@ Abstract.
Recently Rather et al. [9] considered the generalized derivative and the generalized polar derivative and studied the relative position of zeros of generalized derivative and generalized polar derivative with respect to the zeros of polynomial.
In this paper, we establish some inequalities that estimate the maximum modulus of generalized derivative and the generalized polar derivative of the polynomial $P(z)$, which is also the extension of recently known results.
@@ 1. Introduction
Now for the estimation of the lower bounds of $Re\Big(\frac{zP^{\prime}(z)}{P(z)}\Big)$ on $|z|=1$, Dubinin [5] proved the following.
@@
If $P(z)\in P_{n}$ has all its zeros in $|z|\leq 1$, then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
@@
If $P(z)\in P_{n}$ has all its zeros in $|z|\leq k$, $k\leq 1$ then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
@@
As an extension of the inequality $(1)$ to the polar derivative, Aziz [2] proved the following result.
@@
If all the zeros of $P(z)$ lie $|z|\leq k,k\leq 1$, then for $\alpha\in\mathbb{C}$ with $\left|\alpha\right|\geq k,$
@@
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then
@@
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
@@
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq 1$, then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
@@ 2. Main Results
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
@@
If we take $\gamma=(1,1,\cdots,1)$ in Theorem 11, we get the following result.
@@
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
@@
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$
@@
If we take $\gamma=(1,1,\cdots,1)$ in Theorem 22, we get the following result.
@@
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|D_{\alpha}[P](z)|\geq n\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[1+\frac{1}{n}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
For the proof of the theorems we need following lemmas, the first is due to Rather et al. [12].
@@
From a simple consequence of Maximum Modulus theorem, we can get the following result.
@@
[3]. If $P(z)\in\mathcal{P}_{n},$ having all its zeros in the disk $|z|\leq k$ where $k\geq 1,$ then
@@ 3. Proof of Theorems
Since $P(z)$ is a polynomial of degree $n$ has all its zeros in $|z|\leq k$, $k\geq 1$.
@@
$\geq$ $\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{1-\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}{1+\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}\Big)\Big]|G(z)|$
@@
$\text{and so}\displaystyle\max_{|z|=1}\Big|P(kz)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-\frac{z_{j}}{k}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(kz)|$
Now from Lemma 33, we have
$k\displaystyle\max_{|z|=k}\Big|P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|$
$Thus\qquad k\displaystyle\max_{|z|=k}\Big|P^{\gamma}(z)\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|$
$so\displaystyle\max_{|z|=1}\Big|P^{\gamma}(z)\Big|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(z)|$
@@
Since the polynomial $P(z)$ has all its zeros in $|z|\leq k,$ where $k\geq 1.$
Hence the polynomial $G(z)=P(kz)$ has all its zeros in $|z|\leq 1$.
@@
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[G](z)\bigg|\geq$ $\frac{\frac{|\alpha|}{k}-1}{2}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|$
$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|.$
@@
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(kz)|$
$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
With the help of Lemma 33, we have
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
$\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\underset{|z|=1}{max}|P(z)|.$
@@
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$ (14)
@@
On using Lemma 22, this gives
@@
Combining the inequalities $(14)$ and $(15)$, we get
$k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$
@@
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$

### v2

@@ Abstract.
Recently Rather et al. [9] considered the generalized derivative and the generalized polar derivative and studied the relative position of zeros of generalized derivative and generalized polar derivative with respect to the zeros of polynomial. In this paper, we establish some inequalities that estimate the maximum modulus of generalized derivative and the generalized polar derivative of the polynomial $P(z)$, which is also the extension of recently known results.
@@ 1. Introduction
The estimation of the lower bounds of $Re\Big(\frac{zP^{\prime}(z)}{P(z)}\Big)$ on $|z|=1$, Dubinin [5] proved the following.
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ has all its zeros in $|z|\leq 1$, then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ has all its zeros in $|z|\leq k$, $k\leq 1$, then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
@@
As an extension of the inequality $(1)$ to the polar derivative, Aziz [2] proved that
@@
If all the zeros of $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ lie $|z|\leq k,k\leq 1$, then for $\alpha\in\mathbb{C}$ with $\left|\alpha\right|\geq k,$
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq 1$, then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
@@ 2. Main Results
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
@@
If we take $\gamma=(1,1,\cdots,1)$ in Theorem $1$, we get the following result.
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$ (13)
@@
Remark 1.
If we divide inequality (13) by $|\alpha|$ and then letting $|\alpha|\rightarrow\infty$, we get the inequality
$\underset{|z|=1}{max}|P^{\gamma}(z)|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$ (14)
If we take $\gamma=(1,1,\cdots,1)$ in Theorem $2$, we get the following result.
@@
If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|D_{\alpha}[P](z)|\geq n\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[1+\frac{1}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
For the proof of the theorems we need following lemmas. The first is due to Rather et al. [12].
@@
By a simple consequence of Maximum Modulus theorem, we can get the following result.
@@
[3]. If $P(z)$ is a polynomial having all its zeros in the disk $|z|\leq k$ where $k\geq 1,$ then
@@ 3. Proof of Theorems
Since $P(z)$ is a polynomial of degree $n$ having all its zeros in $|z|\leq k$, $k\geq 1$.
@@
Using Lemma 1, we get,
$|zG^{\gamma}(z)|\geq$ $\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{1-\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}{1+\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}\Big)\Big]|G(z)|$
@@
Replacing $G(z)$ by $P(kz)$, we get,
$\displaystyle\max_{|z|=1}\Big|P(kz)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-\frac{z_{j}}{k}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(kz)|$
$k\displaystyle\max_{|z|=k}\Big|P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=k}|P(z)|$
Now from Lemma $3$, we have
$k\displaystyle\max_{|z|=k}\Big|P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|.$
Thus we have,
$k\displaystyle\max_{|z|=k}\Big|P^{\gamma}(z)\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|.$
Using Lemma 2, we get,
$k^{n}\displaystyle\max_{|z|=1}\Big|P^{\gamma}(z)\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|.$
That is,
$\displaystyle\max_{|z|=1}\Big|P^{\gamma}(z)\Big|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(z)|$
@@
Since the polynomial $P(z)$ has all its zeros in $|z|\leq k$ where $k\geq 1.$
Therefore the polynomial $G(z)=P(kz)$ has all its zeros in $|z|\leq 1$.
@@
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[G](z)\bigg|\geq$ $\frac{\frac{|\alpha|}{k}-1}{2}\Big[\wedge+\gamma_{m}\Big(\frac{|k^{n}a_{n}|-|a_{0}|}{|k^{n}a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|$
$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|.$
@@
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(kz)|$
$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
With the help of Lemma $3$, we have
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
$\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\underset{|z|=1}{max}|P(z)|.$
@@
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$ (16)
@@
On using Lemma $2$, this gives
@@
Combining inequalities $(14)$ and $(15)$, we get
$k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$
@@
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$